# SounDiT: Geo-Contextual Soundscape-to-Landscape Generation

### v1

@@ Abstract
We present a novel and practically significant problem—Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation—which aims to synthesize geographically realistic landscape images from environmental soundscapes. Prior audio-to-image generation methods typically rely on general-purpose datasets and overlook geographic and environmental contexts, resulting in unrealistic images that are misaligned with real-world environmental settings. To address this limitation, we introduce a novel geo-contextual computational framework that explicitly integrates geographic knowledge into multimodal generative modeling. We construct two large-scale geo-contextual multimodal datasets, SoundingSVI and SonicUrban, pairing diverse soundscapes with real-world landscape images. We propose SounDiT, a novel Diffusion Transformer (DiT)-based model that incorporates geo-contextual scene conditioning to synthesize geographically coherent landscape images. Furthermore, we propose a practically-informed geo-contextual evaluation framework, the Place Similarity Score (PSS), across element-, scene-, and human perception-levels to measure consistency between input soundscapes and generated landscape images. Extensive experiments demonstrate that SounDiT outperforms existing baselines in both visual fidelity and geographic settings. Our work not only establishes foundational benchmarks for GeoS2L generation but also highlights the importance of incorporating geographic domain knowledge in advancing multimodal generative models, opening new directions at the intersection of generative AI, geography, urban planning, and environmental sciences.
@@ 1 Introduction
“The earth has music for those who listen.” – William Shakespeare
Each place is characterized by a unique set of visual and sounding environments [66, 4, 6, 2], such as natural bird calls in a quiet park starkly contrasting with bustling traffic in downtowns at peak hours. Understanding such geo-contextual linkages regarding how visual and sounding environments co-occur benefits broad applications in geography [74, 14, 39], environmental psychology [53, 26], ecology [60], and urban planning [15]. Integrating these modalities not only advances our knowledge of human-environment interactions [76], but also supports real-world practices [33]. For instance, understanding how the environment shapes human multi-sensory experiences and perceptions [75, 86], informing noise mitigation and place-making practices to benefit human well-being [1, 32], and enhancing inclusive navigation for visually impaired populations [37, 8].
@@
Prior research in the Audio-to-Image (audio2image) stream that translates auditory information into visual representations has offered promising opportunities [70, 61, 80, 79, 5]. However, we observed two noticeable gaps when applying prior studies in geographic contexts and planning practices. First, most audio2image studies rely on generic audio-visual datasets such as object audio [57, 85, 9], human voice [54, 49], weathers [42], and limited scene types [41], and overlook geo-contextual insights such as parks vs. beaches and urban vs. rural settings [81] (See Appendix LABEL:appendix:dataset). Geographers are primarily interested in the environmental context of a soundscape rather than identifying sound sources such as objects or animals. Second, common evaluation metrics employed in prior studies, such as the Fréchet Inception Distance (FID) [30], may fall short in capturing geographic semantic alignment between the input soundscape and generated images. Ideally, the evaluation metrics should consider environmental characteristics and geographic contexts: the place setting depicted in the input soundscape and those represented in the generated images should be similar to ensure geographic contextual coherence and relevance. Furthermore, Diffusion Transformer (DiT) [59] techniques have recently demonstrated impressive results across various tasks such as image [44, 10, 11] and video generations [51, 77, 38, 48]. However, they remain unexplored in audio2image generation and have yet to incorporate geographic knowledge to support applications in geography and urban planning.
@@
Given these challenges, we introduce and formalize a new problem, namely, Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation (Figure 1), to synthesize realistic landscape images from environmental soundscapes, optionally conditioned on scene context. The GeoS2L problem extends traditional audio2image synthesis by incorporating geo-contextual insights with significant practical implications. Here, landscape refers to geographical environments shaped by natural and built features [52]. A soundscape denotes the acoustic environment as perceived by humans at places, comprising sounds from both natural and anthropogenic sources (e.g., bird calls and vehicle noises) [68]. A scene indicates place types such as beach, forest, street, to guide image generation. Our contributions span three key areas: First, we introduce two large-scale, multi-modal geo-contextual datasets: SoundingSVI and SonicUrban, composed of over 330K and 250K soundscape-landscape image pairs. These datasets cover diverse geographic locations and scenes, supporting geo-contextual realistic synthesis. Second, we develop an innovative generative model, SounDiT, which integrates explicit geographic contexts into DiT to effectively translate environmental soundscapes into corresponding geographically coherent landscape images. Third, we design a novel geo-contextual evaluation framework, namely, Place Similarity Scores (PSS), which includes three metrics: element-level, scene-level, and perception-level similarity to assess geographic contexts and environmental consistency of generated landscape images. Our proposed SounDiT achieves state-of-the-art performance on SoundingSVI and SonicUrban. Beyond technical contributions, this work supports applications in geography, environment, and urban planning with significant practical impacts.
@@ 2 Related Work
Audio-to-Image Synthesis aims to translate acoustic features into coherent and meaningful images and has gained increasing attention with the advancement of various multi-modal alignment foundation models [79, 70] and generative techniques [24, 65]. Current audio2image models are broadly divided into GAN-based (e.g., sound2scene [70]) and Diffusion-based (e.g., AudioToken [80], and GluenGen [61]). These models mainly rely on generic datasets which lack geo-contextual information, often limited to object audios [57], human voice [54], weather [42], and broad scene types [41]. As a result, they often produce stylized or unrealistic outputs and fail to capture the geographic coherence between soundscapes and landscapes (See Appendix LABEL:appendix:dataset).
Landscapes and Soundscapes represent two dimensions for characterizing places [46, 19]. Landscapes offer visual representations of the physical layout and built environments, and are commonly captured via street view imagery (SVI) and remote sensing (RS) images [41, 42]. In parallel, soundscapes refer to the acoustic environment perceived at specific places. Existing datasets, such as BirdCLEF [34], BirdSet [64], SoundingEarth [28]), AudioSet [20], UrbanSound8K [67], SONYC [7], EnigenScape [25], and EMO-Soundscape [16], have been utilized to support tasks including monitoring biodiversity [63], noise pollution [32], and human perception analysis [18]. However, prior audio2image studies have overlooked such geographic and environmental insights, limiting their relevance for real-world applications. To address this gap, our work integrates geo-contextual information for multimodal generative modeling by constructing datasets and evaluation frameworks.
3 Method
@@ 3.1 Preliminary
Diffusion Model. We adopt a latent-based diffusion model, consisting of a forward diffusion and backward denoising process conducted in the latent space $\mathbf{z}$. The forward diffusion $\mathbf{q}$ starts from the initial latent $\mathbf{z}_{0}$, gradually adding Gaussian noise over timestep $T$ based on a fixed variance $\alpha$ and defined as:
$q(\mathbf{z}_{t}\mid\mathbf{z}_{t-1})=\mathcal{N}\bigl{(}\mathbf{z}_{t};\sqrt{\alpha_{t}}\,\mathbf{z}_{t-1},(1-\alpha_{t})\,\mathbf{I}\bigr{)},\quad t=1,\dots,T.$
@@
$p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t})=\mathcal{N}\bigl{(}\mathbf{z}_{t-1};\,\mu_{\theta}(\mathbf{z}_{t},t,v),\,\Sigma_{\theta}(\mathbf{z}_{t},t,v)\bigr{)},$
@@
$\hat{\boldsymbol{\epsilon}}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)+s\bigl{[}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,v)-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)\bigr{]}.$
3.2 Geo-Contextual Soundscape-to-Landscape Generation (GeoS2L) Problem Definition
In this work, we formulate our task as the Geo-Contextual Soundscape-to-Landscape (GeoS2L) problem, a specialized variant of the audio2image problem, which aims to generate landscape images from environmental soundscapes guided by an optional scene prompt. We define a landscape $l$ as the geographical and ecological environment shaped by natural and built features, and a soundscape $s$ as the human-perceived acoustic environment at a given place, comprising both natural and anthropogenic sounds such as bird calls and traffic noise. To better support real-world practices, we introduce scene prompt $c$ as an optional semantic label (e.g., park, beach, street) to provide additional geographic semantic context.
Given a dataset $D=\{\,(s_{i},c_{i},l_{i})\mid i=1,\dots,N\},$ our goal is to learn a generator $G\colon(s_{i},c_{i})\mapsto\hat{l}_{i}$ such that the synthesized image $\hat{l}_{i}$ is both visually realistic and geo-contextually consistent to the ground truth $l_{i}$, quantified by relevance function $R(s,c,l)$. This can be formulated as:
$\mathcal{L}\;=\;\mathbb{E}_{(s_{i},c_{i},l_{i})\sim D}\Bigl{[}\,R(s_{i},c_{i},l_{i})\;-\;R\bigl{(}s_{i},c_{i},\hat{l}_{i}\bigr{)}\Bigr{]}.$
Minimizing $\mathcal{L}$ enables the model to capture cross-modal correlations among sounding, semantic and visual features, thereby producing landscape images that reflect both the acoustic characteristics of soundscape and intended scene.
Figure 2: Framework of SounDiT. To effectively capture geographic contexts, SounDiT integrates soundscape, scene, and landscape embeddings from pre-trained encoders as inputs. The landscape embedding is fused with the timestep embedding and fed into a self-attention layer to extract visual features. The scene embedding then guides conditioning via a cross-attention layer. The TA-PMoE module further combines these features with soundscape features. During inference, a geo-contextual RAG module provides additional geographic contexts to enhance alignment with real-world environmental settings.

### v2

# arXiv:2505.12734v2[cs.SD] 27 Feb 2026

@@ Abstract
Recent audio-to-image models have shown impressive performance in generating images of specific objects conditioned on their corresponding sounds. However, these models fail to reconstruct real-world landscapes conditioned on environmental soundscapes. To address this gap, we present Geo-contextual Soundscape-to-Landscape (GeoS2L) generation, a novel and practically significant task that aims to synthesize geographically realistic landscape images from environmental soundscapes. To support this task, we construct two large-scale geo-contextual multi-modal datasets, SoundingSVI and SonicUrban, which pair diverse environmental soundscapes with real-world landscape images. We propose SounDiT, a diffusion transformer (DiT)-based model that incorporates environmental soundscapes and geo-contextual scene conditioning to synthesize geographically coherent landscape images. Furthermore, we propose the Place Similarity Score (PSS), a practically-informed geo-contextual evaluation framework to measure consistency between input soundscapes and generated landscape images. Extensive experiments demonstrate that SounDiT outperforms existing baselines in the GeoS2L, while the PSS effectively captures multi-level generation consistency across element, scene,and human perception. Project page: https://gisense.github.io/SounDiT-Page/
@@ 1 Introduction
“The earth has music for those who listen.”
— William Shakespeare
Environmental soundscapes encode rich geographic information, often in greater detail than the specific object acting as the sound source [67, 4, 6, 1]. A burst of birdsong, for example, implies a leafy trail or an urban green space, while the sound of traffic may evoke a busy streetscape, beyond a single car. Recognizing this strong geo-contextual connection between auditory soundscape and visual landscape, researchers in geography [75, 16, 43], urban planning [17, 39, 31], and environmental psychology [56, 28], have long studied the impacts of soundscapes. For example, how analyzing specific sound types can inform and improve urban design[36], how to mitigate noise through environmental design [7, 8], or which soundscapes are associated with human restoration and well-being [84, 76]. Previous practices in these fields have often been limited to descriptive statistics of the soundscape environment (e.g., decibel levels, sound types) [37, 15], which are not always intuitive and fail to capture the co-located visual characteristics of the scene. The recent emergence of deep generative models for Audio-to-Image (A2I) synthesis [72, 63, 81, 80, 5], which translates auditory information directly into visual representations, offers a promising opportunity to model the geo-contextual soundscape-landscape relationships, unlocking significant practical applications.
@@
However, we observe three noticeable gaps when applying prior A2I in geographic contexts and planning practices for associating soundscapes and landscapes. First, most A2I studies rely on generic audio-visual datasets such as object audio [60, 87, 10], human voice [57, 52], weathers [46], and limited scene types [45]. Consequently, these models link an audio signal to its source (e.g., generating an image of a bird for a bird call) rather than its environment, overlooking geo-contextual insights vital for real-world applications, such as parks vs. beaches and urban vs. rural settings [82]. For geographers and urban planners, the primary interest is not the object making the sound (that bird, that car), but the environmental setting that informs environmental design and noise mitigation strategies. Despite explorations of diffusion-based models in geographic audio-to-image settings [89, 88], these methods still take soundscapes as the sole input and output landscape images, without geographic context. Second, Diffusion Transformer (DiT) [62] techniques have recently demonstrated impressive results across various tasks such as image [48, 12, 11] and video generations [54, 77, 42, 51]. However, they remain unexplored in A2I generation and have yet to incorporate geographic knowledge to support applications in geography and urban planning. Furthermore, common evaluation metrics employed in prior studies, such as the Fréchet Inception Distance (FID) [33], may fall short in capturing geographic semantic alignment between the input soundscape and generated images. Ideally, the evaluation metrics should consider environmental characteristics and geographic contexts: the place setting depicted in the input soundscape and those represented in the generated images should be similar to ensure geographic contextual coherence and relevance.
@@
Given these challenges, we introduce and formalize a new, practically-motivated task: Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation (Figure SounDiT: Geo-Contextual Soundscape-to-Landscape Generation), to synthesize realistic landscape images from environmental soundscapes, optionally conditioned on scene context. The GeoS2L problem extends traditional A2I synthesis by incorporating geo-contextual insights with significant practical implications. Here, landscape refers to geographical environments shaped by natural and built features [55]. A soundscape denotes the acoustic environment as perceived by humans at places, comprising sounds from both natural and anthropogenic sources (e.g., bird calls and vehicle noises) [69]. A scene indicates place types such as beach, forest, street, to guide image generation. Our contributions span three key areas: First, we introduce two large-scale, multi-modal geo-contextual datasets: SoundingSVI and SonicUrban, composed of over 169K and 237K soundscape-landscape image pairs. These datasets cover diverse geographic locations and scenes, supporting geo-contextual realistic synthesis. Second, we develop SounDiT, an innovative generative model that support GeoS2L problem. It introduces three novel, efficient components to integrate geo-contextual information: a MoE Soundscape Conditioning module, a Scene Low-Rank Content Mixer (SLRCM), and a Scene AdaLn (S-AdaLn). These modules effectively inject geographic contexts to environmental soundscapes, enabling the generation of geographically coherent landscape images. Third, we design a novel practically-motivated geo-contextual evaluation framework: Place Similarity Scores (PSS). PSS includes three metrics: element-level, scene-level, and perception-level similarity to assess geographic contexts and environmental consistency of generated landscape images. Extensive experiments demonstrate that our model achieves state-of-the-art performance and delivers strong utility on downstream applications like geography, environment, and urban planning with significant practical impacts. Moreover, our datasets and evaluation metric provide a robust benchmark for the GeoS2L task.
@@ 2 Related Work
Audio-to-Image Synthesis aims to translate acoustic features into coherent and meaningful images and has gained increasing attention with the advancement of various multi-modal alignment foundation models [80, 72] and generative techniques [25, 66]. Current A2I models are broadly divided into GAN-based (e.g., sound2scene [72]) and Diffusion-based (e.g., AudioToken [81], and GluenGen [63]). These models mainly rely on generic datasets which lack geo-contextual information, often limited to object audios [60], human voice [57], weather [46], and broad scene types [45]. As a result, they often produce stylized or unrealistic outputs and fail to capture the geographic coherence between soundscapes and landscapes.
Landscapes and Soundscapes represent two dimensions for characterizing places [50, 20]. Landscapes offer visual representations of the physical layout and built environments, and are commonly captured via street view imagery (SVI) and remote sensing (RS) images [45, 46]. In parallel, soundscapes refer to the acoustic environment perceived at specific places. Existing datasets, such as BirdCLEF [38], BirdSet [65], SoundingEarth [30]), AudioSet [22], UrbanSound8K [68], SONYC [9], EnigenScape [27], and EMO-Soundscape [18], have been utilized to support real-world practices including monitoring biodiversity [64], noise mitigation [35], and human perception analysis [19]. However, prior A2I studies have overlooked such geographic and environmental insights, limiting their real-world applications. To address this gap, our work integrates geo-contextual information for multimodal generative modeling by constructing datasets and evaluation frameworks.
Figure 2: The SounDiT framework. SounDiT encodes soundscape and scene inputs into a shared space with pre-trained encoders, and performs denoising in latent space with DiT blocks equipped with AdaLN-Zero timestep conditioning. Scene information is injected through a Scene Low-Rank Mixer (SLRCM) and a lightweight Scene AdaLN (S-AdaLN) path to enhance geo-contextual consistency, while a Mixture-of-Experts (MoE) Soundscape conditioning module aligns soundscape features with visual landscape tokens.
3 Methodology
@@ 3.1 Preliminary
Diffusion Model. We adopt a latent-based diffusion model, consisting of a forward diffusion and backward denoising process in the latent space $\mathbf{z}$. The forward diffusion $\mathbf{q}$ starts from the initial latent $\mathbf{z}_{0}$, gradually adding Gaussian noise over timestep $T$ based on a fixed variance $\alpha$ and defined as:
$q(\mathbf{z}_{t}\mid\mathbf{z}_{t-1})=\mathcal{N}\bigl(\mathbf{z}_{t};\sqrt{\alpha_{t}}\,\mathbf{z}_{t-1},(1-\alpha_{t})\,\mathbf{I}\bigr).$ (1)
@@
$p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t})=\mathcal{N}\bigl(\mathbf{z}_{t-1};\,\mu_{\theta}(\mathbf{z}_{t},t,v),\,\Sigma_{\theta}(\mathbf{z}_{t},t,v)\bigr).$ (2)
@@
$\hat{\boldsymbol{\epsilon}}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)+s\bigl[\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,v)-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)\bigr].$ (3)
3.2 GeoS2L Task Formulation
We formulate our task as Geo-Contextual Soundscape-to-Landscape (GeoS2L), extending A2I synthesis with geographic contexts. GeoS2L aims to synthesize landscape images from environmental soundscapes guided by an optional scene prompt. We define a landscape $l$ as the geographical–ecological environment shaped by natural and built features, a soundscape $s$ as the human-perceived acoustic environment at a place, including natural and anthropogenic sources (e.g., bird calls and traffic noise). However, real-world soundscape recordings often lack sufficient information to fully capture the visual characteristics of the surrounding environment. For example, a bird call may occur in either rural or urban parks. To provide greater flexibility and user control, we incorporate an optional scene prompt $c$ as the semantic geographic context (e.g., park, beach, street). This prompt serves as an additional geo-contextual conditioning signal to facilitate generative process toward a specific, context-aware landscape. Given a dataset $D=\{(s_{i},c_{i},l_{i})\mid i=1,\dots,N\},$ our goal is to learn a generator $G:\,(s_{i},c_{i})\mapsto\hat{l}_{i}$ such that the synthesized image $\hat{l}_{i}$ is both visually realistic and geo-contextually consistent with the ground truth $l_{i}$, quantified by the relevance function $R(s,c,l)$. This can formulated as
$\mathcal{L}\;=\;\mathbb{E}_{(s_{i},c_{i},l_{i})\sim D}\!\left[\,R(s_{i},c_{i},l_{i})\;-\;R\bigl(s_{i},c_{i},\hat{l}_{i}\bigr)\right].$ (4)
This enables the model to capture cross-modal correlations among acoustic, semantic, and visual cues, producing landscape images that reflect both the soundscape and the intended scene context.